# Gapless symmetry protected topological phases and generalized deconfined critical points from gauging a finite subgroup

### v1

@@ Abstract
Gauging a finite subgroup of a global symmetry can map conventional phases and phase transitions to unconventional ones. In this work, we study, as a concrete example, an emergent $\mathbb{Z}_{2}$-gauged system with global symmetry $U(1)$, namely, the $\mathbb{Z}_{2}$-gauged Bose-Hubbard model both in 1-D and in 2-D. In certain limits, there is an emergent mixed ’t Hooft anomaly between the quotient $\tilde{U}(1)$ symmetry and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry. In 1-D, the superfluid phase is mapped to an intrinsically gapless symmetry-protected topological (SPT) phase, as supported by density-matrix renormalization group (DMRG) calculations. In 2-D, the original superfluid-insulator transition becomes a generalized deconfined quantum critical point (DQCP) between a gapless SPT phase, where a SPT order coexists with Goldstone modes, and a $\tilde{U}(1)$-symmetry-enriched topological (SET) phase. We also discuss the stability of these phases and the critical points to small perturbations and their potential experimental realizations. Our work demonstrates that partial gauging is a simple and yet powerful approach in constructing novel phases and quantum criticalities.
@@ I Introduction
It was realized recently that some DQCPs can be explained using mixed ’t Hooft anomalies, which can be emergent at low energy between the two associated symmetries [3, 4]. The concept of ’t Hooft anomalies, widely studied in high energy physics, has also found deep and broad applications in condensed matter physics since the discovery of topological insulators or, more generally, symmetry protected topological (SPT) phases [5, 6, 7, 8, 9, 10, 11, 12]. These anomalies characterize global symmetries that cannot be gauged consistently. Related to this work, more recently emergent anomalies have been used to construct gapless SPT phases [13, 14, 15, 16, 17] which are “intrinsic” in the sense that not only are the topological edge modes robust against the gapless bulk of the system, but also the SPT nature relies crucially on the gaplessness [18, 19, 20, 21]. ’t Hooft anomalies thus play an important role in extending the Landau paradigm.
@@
Another perspective in extending the Landau paradigm comes from recent development in expanding the definition of “symmetries” to generalized symmetries [22, 23, 24] (non-invertible symmetries included [25, 26]) after it was realized that symmetry generators are essentially topological defects. In particular, ordinary (0-form) symmetries, whose charged objects are 0-dimensional, have been generalized to $p$-form symmetries, whose charged objects are $p$-dimensional. Topologically ordered phases can be interpreted as SSB of some higher-form symmetries. Moreover, it was realized that the Higgs phase can be viewed as a SPT phase protected by higher-form symmetries and is stable to weak explicit breaking of these higher-form symmetries [27, 28].
@@
One more perspective comes from gauging, i.e. coupling systems to dynamical gauge fields. Gauging a theory of matter fields can yield a rich phase diagram. A prominent example is the Fradkin and Shenker model whose phase diagram can contain a confined phase, a Higgs phase, and a deconfined phase [29]. The gauging technique can also be used to extract information in the original system. For example, gauging different SPT phases can lead to distinct topologically ordered phases where quasiparticles have different braiding statistics [30].
@@
Anomalies, higher-form symmetries, and gauging form a powerful toolkit and have led to many interesting new discoveries. It is known that coupling a system to a flat gauge field produces a dual higher-form symmetry and that partially gauging a discrete symmetry can produce a mixed anomaly between the quotient symmetry and the new dual symmetry [31, 32]. It was emphasized in Ref. [33] that gauging a finite subgroup is a general approach to construct exotic critical points from ordinary continuous ones. In 1-D, the critical point where the global symmetry is spontaneously broken is mapped to a DQCP between two SSB phases associated with the quotient symmetry and the dual symmetry [34]. In higher dimensions, it is a generalized DQCP between an ordinary SSB phase and a symmetry enriched topological (SET) phase. We will analyze the generalized DQCP after partial gauging using the new perspectives from higher-form symmetries and mixed anomalies.
@@
In this work, we study the emergent $\mathbb{Z}_{2}$-gauging of a system with global $U(1)$ symmetry. In next section, we describe the general ideas. Starting from Sec. III, we will focus on a concrete model, i.e. the Bose-Hubbard model both in 1-D and in 2-D, coupled to Ising spins on the bonds, where the bosonic parity is effectively gauged. We adapt the argument in the recently proposed “Higgs = SPT” paradigm [27, 28] to argue that the gauged 1-D superfluid phase is actually an intrinsically gapless SPT phase by considering both the periodic boundary condition (PBC) and the open boundary condition (OBC). The critical low energy theory is a $\mathbb{Z}_{2}$-gauged compact boson conformal field theory (CFT). These statements are corroborated by density-matrix renormalization group (DMRG) computations. In Sec. IV, we will argue that in 2-D, the superfluid is also a type of gapless SPT where the gaplessness comes from the Goldstone modes and thus the generalized DQCP is between a gapless SPT phase and a SET phase. We also discuss the effect of some perturbations that explicitly break the dual symmetry, and comment on potential realizations in experiments. We conclude our discussion in Sec. V with some future directions. Some details are presented in the appendices.
@@ II General ideas
Gauging a finite Abelian ordinary (0-form) symmetry in $d$-D space induces a dual $(d-1)$-form symmetry generated by the Wilson operators [22]. The charged objects of the dual $(d-1)$-form symmetry are $(d-1)$-dimensional. One can gauge a finite Abelian normal subgroup $\Gamma$ of the global symmetry $G$ (discrete or continuous), then the global symmetry becomes $G/\Gamma\times\hat{\Gamma}^{(d-1)}$ where $\hat{\Gamma}^{(d-1)}=\hom(\Gamma,U(1))$, the Pontryagin dual of $\Gamma$, is the dual $(d-1)$-form symmetry. If $G$ is a nontrivial extension of $G/\Gamma$ by $\Gamma$, then there is a mixed anomaly between the $G/\Gamma$ and $\hat{\Gamma}^{(d-1)}$[31]. As a corollary, there is no trivially gapped (i.e. nondegenerate, gapped, and symmetric under both symmetries) ground state.
Starting with a general ordinary second order phase transition of Landau type in $d$-D where the global symmetry $G$ is completely spontaneously broken, we can obtain a generalized DQCP by gauging a finite normal subgroup $\Gamma$ of $G$ [33]. The two phases separated by the generalized DQCP are associated with the SSB of $G/\Gamma$ and $\hat{\Gamma}^{(d-1)}$, respectively. In particular, the SSB of a higher-form symmetry $\hat{\Gamma}^{(d-1)}$ ($d\geq 2)$ leads to a topologically order phase [22, 23]. For example, we can gauge the $\Gamma=\mathbb{Z}_{2}$ subgroup of a $\mathbb{Z}_{4}$ clock model in $2$-D where there is an ordinary second order phase transition across which the unbroken $G=\mathbb{Z}_{4}$ is completely broken. The transition point now becomes a generalized DQCP between a SSB phase where the quotient ~$\tilde{\mathbb{Z}}_{2}$ is broken and a SET phase enriched by the quotient ~$\tilde{\mathbb{Z}}_{2}$ (see Ref. [33] and also Appendix A). Using the argument in Refs. [27, 28], we claim that the the quotient ~$\tilde{\mathbb{Z}}_{2}$ SSB phase in fact has boundary modes as long as the dual 1-form ^$\hat{\mathbb{Z}}_{2}$ (as well as the original $\mathbb{Z}_{4}$ symmetry) is preserved. If $G$ is continuous, the SSB of $G/\Gamma$ leads to Goldstone modes, the winding number of which is the charge under the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry. Thus, in the corresponding phase, the boundary modes coexist with the gapless bulk.
@@ III.1 Model
Figure 1: (a) Schematic diagram for the 1-D Bose-Hubbard model (blue sites) coupled to Ising spins (violet bonds). (b) Schematic phase diagram. The original BKT transition between the superfluid and the insulator phase is enriched to a gauged BKT transition between a gapless SPT superfluid phase, protected by $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$, and the insulator phase with $W$ spontaneously broken. (c) Degeneracies in the superfluid and the insulator phase with PBC and OBC, respectively. (d, e) Finite size scaling of the gap $\Delta_{\text{bulk}}$ in both the superfluid ($t=1.0,U=1.0$) and the insulator phase ($t=0.1,U=1.0$). $\Delta_{\text{bulk}}$ is defined to be the gap in spectrum above the (possibly) degenerate ground states. OBC is used in both cases.
@@
Figure 2: (a) Boson pair correlation function, which is shown to follow a power law $\langle b_{i}^{\dagger}b_{i}^{\dagger}b_{j}b_{j}\rangle\sim r^{-\eta_{bb}}$ in the superfluid phase. The inset shows the extrapolation of the exponent $\eta_{bb}$ to the thermodynamic limit using finite-size scaling. Error bars are obtained from the upper and the lower bound of the extrapolation. (b) Gauge-invariant boson correlation, which also follows power law $\langle b_{i}^{\dagger}\sigma^{z}...\sigma^{z}b_{j}\rangle\sim r^{-\eta_{b}}$. The inset shows the extrapolation of $\eta_{b}$. (c) The $\tilde{U}(1)$ disorder operator $|\langle\tilde{X}_{R}(\theta)\rangle|$ decays as power law $r^{-\alpha(\theta)}$ in $r$. The main plot shows the $\theta=2\pi$ case. The inset shows the $\theta$-dependence of the exponent $\alpha$, which is $4\pi$-periodic. $\alpha$ is symmetric about $2\pi$, and a quadratic fit (dashed line) is performed for the segment from 00 to $2\pi$. (d) Subsystem von Neumann entanglement entropy $S_{E}$ as a function of subsystem size $l$. The inset shows the linear dependence of $S_{E}$ on $\lambda(l)\equiv\frac{1}{6}\log\left(\frac{2L}{\pi}\sin\left(\frac{\pi l}{L}\right)\right)$. The central charge $c$, which is given by the slope, is shown to be almost exactly 1. All the main plots are for $L=100$, $t=0.5$ and $U=1.0$.
@@
Before coupling to Ising spins, the Bose-Hubbard model can have two phases: a superfluid or a Mott insulator. The superfluid-insulator transition is a Berezinskii–Kosterlitz–Thouless (BKT) transition point where the transition is due to fluctuations of vortices in the phase [36, 37]. The transition occurs around $t/U\approx 0.3$ [38, 39, 40]. Note that in the supefluid phase, the $U(1)$ symmetry is not broken due to the celebrated Mermin-Wagner theorem [41], but there is a quasi-long range order, where $\langle b^{\dagger}_{i}b_{j}\rangle\sim r^{-\eta_{b}}$ decays algebraically for large $r=|j-i|$ with $\eta_{b}=\tilde{K}/2$, $\tilde{K}$ being the Luttinger parameter. Similarly, the disorder parameter $|\langle X_{R}(\theta)\rangle|=|\langle\prod_{i\in R}e^{i\theta n_{i}}\rangle|$, where $R$ is a line segment with $r=|R|$, decays algebraically.
Typical phase transitions are insensitive to boundary conditions in the thermodynamical limit. Gauging the $\mathbb{Z}_{2}$ subgroup is equivalent to averaging over untwisted and twisted sectors. Thus, gauging the $\mathbb{Z}_{2}$ subgroup does not change the position of critical point in the phase diagram. A continuous phase transition in the gauged model is directly inherited from the ungauged one but with many new features due to the interplay between the quotient $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. The original BKT transition is now gauged (see Fig. 1(b)).
@@
On the other hand, the new Ising degrees of freedom also behave differently in the superfluid and the insulator phase. As a result of the emergent gauge constraints, the relation $\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle=\langle\prod_{i\in R}(-1)^{n_{i}}\rangle=\langle\tilde{X}_{R}(2\pi)\rangle$ holds in either phase. It relates an “order parameter” associated with $W$ to a disorder parameter associated with $\tilde{U}(1)$. As we will discuss soon, this is a manifestation of the emergent mixed anomaly between $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. In the insulator case, the magnetization $|\langle\sigma_{i}^{x}\rangle|$ in the thermodynamical limit is nonzero, $W$ is spontaneously broken, and the spin-spin correlation $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|$ saturates to a constant. In the superfluid phase, the bulk magnetization vanishes and the bulk spin-spin correlation $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|\sim r^{-\eta_{s}}$ decays algebraically regardless of whether PBC or OBC is used. See Fig. 3(a) for the bulk correlation with OBC  [^2] Note that, due to the edge SSB that we will discuss later, the bulk magnetization remains finite, albeit small, for finite systems. In this plot, we have subtracted the expectation $\langle\sigma_{i}^{x}\rangle\langle\sigma_{j}^{x}\rangle$ and limited $r$ to be far away from the edges. .

### v2

# Gapless symmetry-protected topological phases and generalized deconfined critical points from gauging a finite subgroup

@@ Abstract
Gauging a finite subgroup of a global symmetry can map conventional phases and phase transitions to unconventional ones. In this work, we study, as a concrete example, an emergent $\mathbb{Z}_{2}$-gauged system with global symmetry $U(1)$, namely, the $\mathbb{Z}_{2}$-gauged Bose-Hubbard model both in 1-D and in 2-D. In certain limits, there is an emergent mixed ’t Hooft anomaly between the quotient $\tilde{U}(1)$ symmetry and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry. In 1-D, the superfluid phase is mapped to an intrinsically gapless symmetry-protected topological (SPT) phase, as supported by density-matrix renormalization group (DMRG) calculations. In 2-D, the original superfluid-insulator transition becomes a generalized deconfined quantum critical point (DQCP) between a gapless SPT phase, where an SPT order coexists with Goldstone modes, and a $\tilde{U}(1)$-symmetry-enriched topological (SET) phase. We also discuss the stability of these phases and the critical points to small perturbations and their potential experimental realizations. Our work demonstrates that partial gauging is a simple and yet powerful approach in constructing novel phases and quantum criticalities.
@@ I Introduction
It was realized recently that some DQCPs can be explained using mixed ’t Hooft anomalies, which can be emergent at low energy between the two associated symmetries [3, 4]. The concept of ’t Hooft anomalies, widely studied in high energy physics, has also found deep and broad applications in condensed matter physics since the discovery of topological insulators or, more generally, symmetry-protected topological (SPT) phases [5, 6, 7, 8, 9, 10, 11, 12]. These anomalies characterize global symmetries that cannot be gauged consistently. Related to this work, more recently emergent anomalies have been used to construct gapless SPT phases [13, 14, 15, 16, 17] which are “intrinsic” in the sense that not only are the topological edge modes robust against the gapless bulk of the system, but also the SPT nature relies crucially on the gaplessness [18, 19, 20, 21]. ’t Hooft anomalies thus play an important role in extending the Landau paradigm.
@@
Another perspective in extending the Landau paradigm comes from recent development in expanding the definition of “symmetries” to generalized symmetries [22, 23, 24] (noninvertible symmetries included [25, 26]) after it was realized that symmetry generators are essentially topological defects. In particular, ordinary (0-form) symmetries, whose charged objects are 0-dimensional, have been generalized to $p$-form symmetries, whose charged objects are $p$-dimensional. Topologically ordered phases can be interpreted as SSB of some higher-form symmetries. Moreover, it was realized that the Higgs phase can be viewed as an SPT phase protected by higher-form symmetries and is stable to weak explicit breaking of these higher-form symmetries [27, 28].
@@
One more perspective comes from gauging, i.e., coupling systems to dynamical gauge fields. Gauging a theory of matter fields can yield a rich phase diagram. A prominent example is the Fradkin and Shenker model whose phase diagram can contain a confined phase, a Higgs phase, and a deconfined phase [29]. The gauging technique can also be used to extract information in the original system. For example, gauging different SPT phases can lead to distinct topologically ordered phases where quasiparticles have different braiding statistics [30].
@@
Anomalies, higher-form symmetries, and gauging form a powerful toolkit and have led to many interesting new discoveries. It is known that coupling a system to a flat gauge field produces a dual higher-form symmetry and that partially gauging a discrete symmetry can produce a mixed anomaly between the quotient symmetry and the new dual symmetry [31, 32]. It was emphasized in Ref. [33] that gauging a finite subgroup is a general approach to construct exotic critical points from ordinary continuous ones. In 1-D, the critical point where the global symmetry is spontaneously broken is mapped to a DQCP between two SSB phases associated with the quotient symmetry and the dual symmetry [34]. In higher dimensions, it is a generalized DQCP between an ordinary SSB phase and a symmetry-enriched topological (SET) phase. We will analyze the generalized DQCP after partial gauging using the new perspectives from higher-form symmetries and mixed anomalies.
@@
In this work, we study the emergent $\mathbb{Z}_{2}$-gauging of a system with global $U(1)$ symmetry. In next section, we describe the general ideas. Starting from Sec. III, we will focus on a concrete model, i.e., the Bose-Hubbard model both in 1-D and in 2-D, coupled to Ising spins on the bonds, where the bosonic parity is effectively gauged. We adapt the argument in the recently proposed “Higgs = SPT” paradigm [27, 28] to argue that the gauged 1-D superfluid phase is actually an intrinsically gapless SPT phase by considering both the periodic boundary condition (PBC) and the open boundary condition (OBC). The critical low energy theory is a $\mathbb{Z}_{2}$-gauged compact boson conformal field theory (CFT). These statements are corroborated by density-matrix renormalization group (DMRG) computations. In Sec. IV, we will argue that in 2-D, the superfluid is also a type of gapless SPT where the gaplessness comes from the Goldstone modes and thus the generalized DQCP is between a gapless SPT phase and a SET phase. We also discuss the effect of some perturbations that explicitly break the dual symmetry, and comment on potential realizations in experiments. We conclude our discussion in Sec. V with some future directions. Some details are presented in the appendices.
@@ II General ideas
Gauging a finite Abelian ordinary (0-form) symmetry in $d$-D space induces a dual $(d-1)$-form symmetry generated by the Wilson operators [22]. The charged objects of the dual $(d-1)$-form symmetry are $(d-1)$-dimensional. One can gauge a finite Abelian normal subgroup $\Gamma$ of the global symmetry $G$ (discrete or continuous), then the global symmetry becomes $G/\Gamma\times\hat{\Gamma}^{(d-1)}$ where $\hat{\Gamma}^{(d-1)}=\hom(\Gamma,U(1))$, the Pontryagin dual of $\Gamma$, is the dual $(d-1)$-form symmetry. If $G$ is a nontrivial extension of $G/\Gamma$ by $\Gamma$, i.e., $G\ncong G/\Gamma\times\Gamma$, then there is a mixed anomaly between the $G/\Gamma$ and $\hat{\Gamma}^{(d-1)}$[31]. As a corollary, there is no trivially gapped (i.e., nondegenerate, gapped, and symmetric under both symmetries) ground state.
Starting with a general ordinary second order phase transition of Landau type in $d$-D where the global symmetry $G$ is completely spontaneously broken, we can obtain a generalized DQCP by gauging a finite normal subgroup $\Gamma$ of $G$ [33]. The two phases separated by the generalized DQCP are associated with the SSB of $G/\Gamma$ and $\hat{\Gamma}^{(d-1)}$, respectively. In particular, the SSB of a higher-form symmetry $\hat{\Gamma}^{(d-1)}$ ($d\geq 2)$ leads to a topologically order phase [22, 23]. For example, we can gauge the $\Gamma=\mathbb{Z}_{2}$ subgroup of a $\mathbb{Z}_{4}$ clock model in $2$-D where there is an ordinary second order phase transition across which the unbroken $G=\mathbb{Z}_{4}$ is completely broken. The transition point now becomes a generalized DQCP between a SSB phase where the quotient ~$\tilde{\mathbb{Z}}_{2}$ is broken and a SET phase enriched by the quotient ~$\tilde{\mathbb{Z}}_{2}$ (see Ref. [33] and also Appendix A). Using the argument in Refs. [27, 28], we claim that the quotient ~$\tilde{\mathbb{Z}}_{2}$ SSB phase in fact has boundary modes as long as the dual 1-form ^$\hat{\mathbb{Z}}_{2}$ (as well as the original $\mathbb{Z}_{4}$ symmetry) is preserved. If $G$ is continuous, the SSB of $G/\Gamma$ leads to Goldstone modes, the winding number of which is the charge under the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry. Thus, in the corresponding phase, the boundary modes coexist with the gapless bulk.
@@ III.1 Model
Figure 1: (a) Schematic diagram for the 1-D Bose-Hubbard model (blue sites) coupled to Ising spins (violet bonds). (b) Schematic phase diagram. The original BKT transition between the superfluid and the insulator phase is enriched to a gauged BKT transition between a gapless SPT superfluid phase, protected by $U(1)$ and $W$, and the insulator phase with $W$ spontaneously broken. (c) Degeneracies in the superfluid and the insulator phase with PBC and OBC, respectively. (d, e) Finite size scaling of the gap $\Delta_{\text{bulk}}$ in both the superfluid ($t=1.0,U=1.0$) and the insulator phase ($t=0.1,U=1.0$). $\Delta_{\text{bulk}}$ is defined to be the gap in spectrum above the (possibly) degenerate ground states. OBC is used in both cases.
@@
Figure 2: (a) Boson pair correlation function, which is shown to follow a power law $\langle b_{i}^{\dagger}b_{i}^{\dagger}b_{j}b_{j}\rangle\sim r^{-\eta_{bb}}$ in the superfluid phase. The inset shows the extrapolation of the exponent $\eta_{bb}$ to the thermodynamic limit using finite-size scaling. Error bars are obtained from the upper and the lower bound of the extrapolation. (b) Gauge-invariant boson correlation, which also follows power law $\langle b_{i}^{\dagger}\sigma^{z}...\sigma^{z}b_{j}\rangle\sim r^{-\eta_{b}}$. The inset shows the extrapolation of $\eta_{b}$. (c) The $\tilde{U}(1)$ disorder operator $|\langle\tilde{X}_{R}(\theta)\rangle|$ decays as power law $r^{-\alpha(\theta)}$ in $r$. The main plot shows the $\theta=2\pi$ case. The inset shows the $\theta$-dependence of the exponent $\alpha$, which is $4\pi$-periodic. $\alpha$ is symmetric about $2\pi$, and a quadratic fit (dashed line) is performed for the segment from 00 to $2\pi$. (d) Subsystem von Neumann entanglement entropy $S_{E}$ as a function of subsystem size $l$. The inset shows the linear dependence of $S_{E}$ on $\lambda(l)\equiv\frac{1}{6}\ln\left(\frac{2L}{\pi}\sin\left(\frac{\pi l}{L}\right)\right)$. The central charge $c$, which is given by the slope, is shown to be almost exactly 1. Recently, the entanglement spectrum of gapless SPT phases has also been studied and the entanglement spectrum also has degeneracy [36]. All the main plots are for $L=100$, $t=0.5$ and $U=1.0$.
@@
Before coupling to Ising spins, the Bose-Hubbard model can have two phases: a superfluid or a Mott insulator. The superfluid-insulator transition is a Berezinskii–Kosterlitz–Thouless (BKT) transition point where the transition is due to fluctuations of vortices in the phase [37, 38]. The transition occurs around $t/U\approx 0.3$ [39, 40, 41]. Note that in the superfluid phase, the $U(1)$ symmetry is not broken due to the celebrated Mermin-Wagner theorem [42], but there is a quasi-long range order, where $\langle b^{\dagger}_{i}b_{j}\rangle\sim r^{-\eta_{b}}$ decays algebraically for large $r=|j-i|$ with $\eta_{b}=\tilde{K}/2$, $\tilde{K}$ being the Luttinger parameter. Similarly, the disorder parameter $|\langle X_{R}(\theta)\rangle|=|\langle\prod_{i\in R}e^{i\theta n_{i}}\rangle|$, where $R$ is a line segment with $r=|R|$, decays algebraically. Here, the disorder operator $X_{R}(\theta)$, unlike the symmetry generator $X(\theta)$ in Eq. (5), only acts on a subset, $R$, of the total space. If the $U(1)$ symmetry is spontaneously broken, this disorder operator effectively creates two defects at the two ends of $R$. Then the expectation of this operator effectively detects the long-range correlation of the two defects, similar to that in the Ising model [43]. In the 1-D critical superfluid phase, it decays in an algebraic law, similar to the correlation function of two order operators.
Typical phase transitions are insensitive to boundary conditions in the thermodynamical limit. Gauging the $\mathbb{Z}_{2}$ subgroup is equivalent to averaging over untwisted and twisted sectors. Thus, gauging the $\mathbb{Z}_{2}$ subgroup does not change the position of critical point in the phase diagram. A continuous phase transition in the gauged model is directly inherited from the ungauged one but with many new features due to the interplay between the quotient $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. The original BKT transition is now gauged [see Fig. 1(b)].
@@
On the other hand, the new Ising degrees of freedom also behave differently in the superfluid and the insulator phase. As a result of the emergent gauge constraints, the relation $\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle=\langle\prod_{i\in R}(-1)^{n_{i}}\rangle=\langle\tilde{X}_{R}(2\pi)\rangle$ holds in either phase. It relates an “order parameter” associated with $W$ to a disorder parameter associated with $\tilde{U}(1)$. As we will discuss soon, this is a manifestation of the emergent mixed anomaly between $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. In the insulator case, the magnetization $|\langle\sigma_{i}^{x}\rangle|$ in the thermodynamical limit is nonzero, $W$ is spontaneously broken, and the spin-spin correlation $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|$ saturates to a constant. In the superfluid phase, the bulk magnetization vanishes and the bulk spin-spin correlation $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|\sim r^{-\eta_{s}}$ decays algebraically regardless of whether PBC or OBC is used. See Fig. 3(a) for the bulk correlation with OBC  [^2] Note that, due to the edge SSB that we will discuss later, the bulk magnetization remains finite, albeit small, for finite systems. In this plot, we have subtracted the expectation $\langle\sigma_{i}^{x}\rangle\langle\sigma_{j}^{x}\rangle$ and limited $r$ to be far away from the edges. . Later on, we will show that these critical exponents are not all independent. In fact, they are compatible with the predictions from a compact boson CFT.